# JavisDiT [Uncaptioned image]: Joint Audio-Video Diffusion Transformer with Hierarchical Spatio-Temporal Prior Synchronization

### v1

@@ C.2.2 Spatial-Temporal VA-Encoding
For a given video clip with the corresponding audio, we first use OpenSora’s [65] and AudioLDM2’s [27] VAE-encoders to respectively map the raw data to the feature space, and derive a lightweight VA-Fuser module to extract the spatio-temporal embeddings from the video-audio pair. As shown in Fig. A2, the raw video $V\in\mathbb{R}^{T\times(H\times W)\times C}$ will be encoded and reshaped to a video embedding $\mathbf{v}=E_{v}(V)\in\mathbb{R}^{T_{v}\times S_{v}\times D_{v}}$, where the downsampling rate for ($T$,$H$,$W$) is ($4$, $8$, $8$), and the output channel $D=4$. Similarly, the raw audio will be first transformed to the MelSpectrogram $A\in\mathbb{R}^{R\times M\times 1}$, and then encoded to an audio embedding $\mathbf{a}=E_{a}(A)\in\mathbb{R}^{T_{a}\times S_{a}\times D_{a}}$. We use a linear layer to unify video and audio channels into the hidden size $D=512$.
@@ C.2.3 Contrastive Training Losses
    Token-level hinge loss: $\mathcal{L}_{token}(\mathbf{p},\mathbf{e}^{+},\mathbf{e}^{-})=\left|1.0-sim(\mathbf{p},\mathbf{e}^{+})\right|+\left|1.0+sim(\mathbf{p},\mathbf{e}^{-})\right|$, where $sim$ refers to the cosine similarity.
@@
$\mathcal{L}_{\text{contrast}}(\mathbf{p},\mathbf{e}^{+},\mathbf{e}^{-})=$ $\mathcal{L}_{token}(\mathbf{p},\mathbf{e}^{+},\mathbf{e}^{-})+$ (A2)
$\mathcal{L}_{disc}(\mathbf{p},\mathbf{e}^{+},\mathbf{e}^{-})+$
$\mathcal{L}_{vad}(\mathbf{e}^{+},\mathbf{e}^{-})+$
$\mathcal{L}_{reg}(\mathbf{p},\mathbf{e}^{+})\,.$
Specifically, we take the text prompts $s$ and estimated prior $(\mathbf{p}_{s},\mathbf{p}_{t})$ as an anchor, and treat the corresponding synchronous video-audio ($\mathbf{v},\mathbf{a}$) pairs from training datasets as positive samples. Then, we either synthesize an asynchronous video $\mathbf{v}^{-}$ or audio $\mathbf{a}^{-}$, and take the original $\mathbf{a}$ or $\mathbf{v}$ to form negative pairs. It can be formulated as $(\mathbf{e}_{s}^{+},\mathbf{e}_{t}^{+})$=$\mathcal{E}_{\psi}(\mathbf{v},\mathbf{a})$, $(\mathbf{e}_{s}^{-},\mathbf{e}_{t}^{-})$=$\mathcal{E}_{\psi}(\mathbf{v},\mathbf{a}^{-})\;\text{or}\;\mathcal{E}_{\psi}(\mathbf{v}^{-},\mathbf{a})$. The VA-Encoder $\mathcal{E}_{\psi}$ is detailed in Sec. C.2.2, and the negative sample construction is introduced in Sec. C.2.4. We also discuss the efficacy of each loss function in Sec. E.3
@@ C.2.4 Negative Sample Construction
Given a synchronized video-audio pair from the training set, we design two approaches to synthesize easy and hard asynchronous videos/audios for optimization. (1) For easy negatives, we take AudioLDM2 [27] to generate arbitrary audios from the text without referencing the corresponding video, which naturally results in asynchronous video-audio pairs. (2) For hard negatives, various augmentation strategies are employed to add, remove, or modify elements in either the original video or audio from GT-pairs, creating more fine-grained spatio-temporally asynchronous pairs. As detailed below, these methods further enhance the discriminative power of the ST-Prior, and Fig. A3 and Fig. A4 showcase some representative augmentation examples.
@@
  • Adding Subject Trajectories (Fig. A3(B)). The trajectories from the SA-V dataset [39] are overlaid to simulate new sound-producing objects. Trajectories are preprocessed using RepViT-SAM [50] for interpolation from 6fps to 24fps, retaining objects with an average size $>32\times 32$ pixels. During augmentation, a random trajectory is selected from the pool and added to the video.
@@
All audio augmentations (except temporal shifting) focus on sound-producing intervals, avoiding silent or noisy segments. We use QwenPlusAPI [^3] https://help.aliyun.com/zh/model-studio/developer-reference/use-qwen-by-calling-api to extract potential sound-producing objects from audio descriptions, and utilize AudioSep [30] to separate sound sources and labels intervals using RMS. Separation results are precomputed and stored as metadata, guiding augmentations on target audios.
@@
  • Audio Volume Adjustment (Fig. A4(C)). The target audio’s volume is modified using a randomly selected transformation (e.g., cosine, sine, linear increase/decrease) to adjust its amplitude dynamically.
@@
  • Audio Speed Adjustment (Fig. A4(G)). Playback speed within the target interval is altered to 0.5$\times$ or 2$\times$, causing another type of timeline desynchronization.
@@ D.1 Motivation
A strong generative model must produce diverse video content and audio types while ensuring fine-grained spatio-temporal synchronization. However, we have noticed that current benchmarks and metrics lack sufficient diversity and robustness for comprehensive evaluation. For instance, the commonly used AIST++ [22] and Landscape [21] benchmarks focus only on limited scenarios (human dancing for AIST++ and natural scenarios for landscape) with merely 20-100 test samples, and the video-audio synchronization metric AV-Align [60] cannot produce accurate evaluation on complicated scenarios, e.g., with multiple sounding events [17, 31]. On the other hand, although TAVGBench [31] provides 3,000 test samples (but not released yet), it lacks detailed analysis of different scenarios in practical audio-video generations. We are therefore motivated to propose a more challenging benchmark to evaluate generative models from various dimensions on a larger data scale in Sec. 4.1, and also introduce a robust metric to quantify the fine-grained spatio-temporal alignment between generated video and audio pairs in Sec. 4.2. Here we supplement more details of the benchmark construction and metric verification.
@@ D.2 Taxonomy
Urban Scenario Outdoor spaces shaped by human activity, including cities, villages, streets, and parks.
Living Scenario Indoor environments where daily human activities occur, like houses, schools, and shopping malls.
Industrial Scenario Work-oriented spaces related to industrial or energy activities, such as factories, construction sites, and mines.
Virtual Scenario Imaginative or abstract settings, including virtual worlds, sci-fi cities, and artistic installations.
@@
2D-Animate Styles like hand-drawn animation, flat animation, cartoon styles, or watercolor illustrations.
3D-Animate Photorealistic styles, sci-fi/magical effects, CG (Computer Graphics), or steampunk aesthetics.
@@
Biological Sounds Sounds produced by living creatures (e.g. animals, birds). This includes vocalizations such as barking, chirping, growling, as well as non-vocal human sounds like heartbeat, and other physical noises.
Mechanical Sounds Sounds generated by man-made machines, devices, or mechanical processes. This includes the noise of engines, motors, appliances, and any mechanical or electronic noise. This category also includes malfunction sounds (e.g., malfunctioning machinery or alarms).
Musical Sounds Sounds related to music or musical performance, including both human-generated and instrument-generated sounds and melodies. This category covers singing, instrumental performances, as well as background music used in various media formats.
Speech Sounds Sounds generated from human speech, whether in conversation, dialogue, public speeches, debates, interviews, or monologues. This category specifically covers linguistic communication in various contexts, whether formal, informal, or contentious.
@@
Multiple Subject There are multiple primary objects that (or potentially can) make sounds in the scene.
Off-screen Sound The source of the sound is not visible in the scene but logically exists (e.g., a car engine outside the camera view).
@@
Sequential Events There are multiple events occurring sequentially, with no overlap. For example, “the applause begins after the music performance ends.”
Simultaneous Events Multiple audio sources are present simultaneously, such as “a person speaking while music plays in the background.”
To comprehensively evaluate the capabilities of joint video-audio generation models, we designed five evaluation dimensions from different aspects: (1) Event Scenario, (2) Video Style, (3) Sound Type, (4) Spatial Subjects, and (5) Temporal Composition. Leveraging GPT-4 [1], we enumerated 3-5 primary categories under each aspect, contributing to a hierarchical categorization system with 5 dimensions and 19 categories. Tab. A2 presents the detailed definitions and clarifications.
@@ D.3.1 Collection
We collected data from two sources to construct the benchmark. First, we incorporate existing datasets’ test sets, including those in the JAVG domain like Landscape [21] and AIST++ [22] (despite their limited coverage) and others from sounding-video understanding tasks such as FAVDBench [46] and AVSBench [66]. Second, we expanded the benchmark by crawling videos uploaded to YouTube between June 2024 and December 2024 to prevent data leakage in previous methods’ training scope [31]. Using the previously defined categorization system, we prompt GPT4 to generate potential keywords for specific categories, enabling targeted video collection and avoiding indiscriminate gathering of homogeneous data, which significantly improves curation efficiency. This stage yields around 30K sounding video candidates.
@@ D.3.2 Quality-based Pre-Filtering
Table A3: Data source composition before and after a series of filtering strategies of our JavisBench
YouTube FAVDBench AVSBench Landscape AIST++ Total
Before Filtering 30,107 1,000 804 100 20 32,031
After Filtering 8,507 833 680 100 20 10,140
Following OpenSora [65], we utilize a series of filtering tools to improve data quality, including:
@@
    Aesthetic Filtering. We filter out videos with aesthetic scores [45] lower than 4.5, remaining in 70K clips.
@@
    Optical-flow Filtering. We use UniMatch [57] to estimate the motion quality of the given videos, and remove the (static) videos whose score is lower than 0.1. This produces 46K candidates.
@@
    OCR Filtering. We use DBNet [23] to detect and filter out the videos containing more than 5 text regions, resulting in a smaller 30K scope.
@@ D.3.3 Annotation
  • First, we use Qwen2-VL-72B [52] to generate detailed captions for videos. Metadata from the data source (e.g., object labels or YouTube keywords) is included in the context to enhance caption quality.
  • Next, we use Qwen2-Audio-7B [5] to caption the audio in detail. We do not incorporate the previously generated video captions as contextual input, as it produces even more hallucinations with the input bias (i.e., wrongly identifying a sound that will happen in corresponding visual scenarios but actually does not exist in the current audio).
  • Then, we employ Qwen2.5-72B-Instruct [58] to merge video and audio captions into a unified text prompt, serving as the textual condition for the JAVG task. During this process, Qwen2.5-72B is prompted to reason on the video-audio captions and identify apparent logical mistakes (e.g., there is a dog in the video but Qwen2-Audio gives sounds from a car engine) or missing captions (e.g., unknown sound source).
@@ D.3.4 Content-based Post-Filtering
To build a diversified and balanced benchmark to evaluate joint audio-video generation, we conduct another post-filtering based on the categorization results obtained above. In particular, we further remove videos that only contain background music and speech voice. By doing so, nearly half of videos are filtered out, since a large part of YouTube videos rely on music and voice acting to attract viewers. After human checking, we obtained 10,140 samples in our JavisBench with diverse data sources and fine-grained category annotations, setting a new standard to facilitate a comprehensive evaluation for future joint audio-video generation (JAVG) methods. The category statistics can be found in Fig. 4, and Tab. A3 displays the data sources before and after our filtering strategies.
@@ D.4.1 Implementation Details
Given a generated video-audio pair ($V$, $A$), we use ImageBind [11] to estimate audio-visual synchrony with following steps:
@@
    Temporal Segmentation via Sliding Windows. ($V$, $A$) is chunked into several segments with 2-seconds window size and 1.5-seconds overlap ($\mathcal{C}=\{(V_{1},A_{1}),(V_{2},A_{2}),\cdots(V_{W},A_{W})\}$). The 2-second window ensures compatibility with the audio encoder’s minimum processing length [11], preventing suboptimal feature extraction. The 1.5-second overlap enhances continuity and robustness, allowing each frame to be evaluated within multiple temporal contexts. This mitigates artifacts caused by abrupt segmentation boundaries.
@@
    Frame-wise Audio-Visual Similarity Computation. Inspired by Mao et al. [31], we calculate the cosine similarity between each frames (=$\mathcal{F}=\{V_{i,1},V_{i,2},\cdots V_{i,w}\}$) and the whole audio clip $A_{i}$ by using ImageBind’s vision and audio encoders ($E_{v}$, $E_{a}$). Frame-wise similarity captures fine-grained temporal dynamics, ensuring transient events (e.g., lip movements, object interactions) are accurately modeled.
@@
$S_{\textit{Javis}}$ $=\frac{1}{W}\sum_{i=1}^{W}\sigma(V_{i},A_{i}),$ (A3)
$\sigma(V_{i},A_{i})$ $=\frac{1}{k}\sum_{j=1}^{k}\mathop{\text{top-}k}\limits_{\min}\{\cos\left(E_{v}(V_{i,j}),E_{a}(A_{i})\right)\}\,.$

### v2

@@ C.2.2 Spatial-Temporal VA-Encoding
For a given video clip with the corresponding audio, we first use OpenSora’s (Zheng et al., 2024) and AudioLDM2’s (Liu et al., 2024b) VAE-encoders to respectively map the raw data to the feature space, and derive a lightweight VA-Fuser module to extract the spatio-temporal embeddings from the video-audio pair. As shown in Fig. A3, the raw video $V\in\mathbb{R}^{T\times(H\times W)\times C}$ will be encoded and reshaped to a video embedding $\mathbf{v}=E_{v}(V)\in\mathbb{R}^{T_{v}\times S_{v}\times D_{v}}$, where the downsampling rate for ($T$,$H$,$W$) is ($4$, 88, 88), and the output channel $D=4$. Similarly, the raw audio will be first transformed to the MelSpectrogram $A\in\mathbb{R}^{R\times M\times 1}$, and then encoded to an audio embedding $\mathbf{a}=E_{a}(A)\in\mathbb{R}^{T_{a}\times S_{a}\times D_{a}}$. We use a linear layer to unify video and audio channels into the hidden size $D=512$.
@@ C.2.3 Contrastive Training Losses
    Token-level hinge loss: $\mathcal{L}_{token}(\mathbf{p},\mathbf{e}^{+},\mathbf{e}^{-})=\left|1.0-sim(\mathbf{p},\mathbf{e}^{+})\right|+\left|1.0+sim(\mathbf{p},\mathbf{e}^{-})\right|$, where simsim refers to the cosine similarity.
@@
$\mathcal{L}_{\text{contrast}}(\mathbf{p},\mathbf{e}^{+},\mathbf{e}^{-})=$ $\mathcal{L}_{token}(\mathbf{p},\mathbf{e}^{+},\mathbf{e}^{-})+\mathcal{L}_{disc}(\mathbf{p},\mathbf{e}^{+},\mathbf{e}^{-})+\mathcal{L}_{vad}(\mathbf{e}^{+},\mathbf{e}^{-})+\mathcal{L}_{reg}(\mathbf{p},\mathbf{e}^{+})\,.$ (A2)
Specifically, we take the text prompts $s$ and estimated prior $(\mathbf{p}_{s},\mathbf{p}_{t})$ as an anchor, and treat the corresponding synchronous video-audio ($\mathbf{v},\mathbf{a}$) pairs from training datasets as positive samples. Then, we either synthesize an asynchronous video $\mathbf{v}^{-}$ or audio $\mathbf{a}^{-}$, and take the original $\mathbf{a}$ or $\mathbf{v}$ to form negative pairs. It can be formulated as $(\mathbf{e}_{s}^{+},\mathbf{e}_{t}^{+})$$\mathcal{E}_{\psi}(\mathbf{v},\mathbf{a})$, $(\mathbf{e}_{s}^{-},\mathbf{e}_{t}^{-})$$\mathcal{E}_{\psi}(\mathbf{v},\mathbf{a}^{-})\;\text{or}\;\mathcal{E}_{\psi}(\mathbf{v}^{-},\mathbf{a})$. The VA-Encoder $\mathcal{E}_{\psi}$ is detailed in Sec. C.2.2, and the negative sample construction is introduced in Sec. C.2.4. We also discuss the efficacy of each loss function in Sec. E.4
@@ C.2.4 Negative Sample Construction
Given a synchronized video-audio pair from the training set, we design two approaches to synthesize easy and hard asynchronous videos/audios for optimization. (1) For easy negatives, we take AudioLDM2 (Liu et al., 2024b) to generate arbitrary audios from the text without referencing the corresponding video, which naturally results in asynchronous video-audio pairs. (2) For hard negatives, various augmentation strategies are employed to add, remove, or modify elements in either the original video or audio from GT-pairs, creating more fine-grained spatio-temporally asynchronous pairs. As detailed below, these methods further enhance the discriminative power of the ST-Prior, and Fig. A4 and Fig. A5 showcase some representative augmentation examples.
@@
  • Adding Subject Trajectories (Fig. A4(B)). The trajectories from the SA-V dataset (Ravi et al., 2024) are overlaid to simulate new sound-producing objects. Trajectories are preprocessed using RepViT-SAM (Wang et al., 2024a) for interpolation from 6fps to 24fps, retaining objects with an average size $32\times 32$ pixels. During augmentation, a random trajectory is selected from the pool and added to the video.
@@
All audio augmentations (except temporal shifting) focus on sound-producing intervals, avoiding silent or noisy segments. We use QwenPlusAPI [^3] https://help.aliyun.com/zh/model-studio/developer-reference/use-qwen-by-calling-api to extract potential sound-producing objects from audio descriptions, and utilize AudioSep (Liu et al., 2024c) to separate sound sources and labels intervals using RMS. Separation results are precomputed and stored as metadata, guiding target augmentations.
@@
  • Audio Volume Adjustment (Fig. A5(C)). The target audio’s volume is modified by a randomly selected transformation (e.g., cosine, sine, linear increase/decrease) to adjust its amplitude.
@@
[…]
@@ D.1 Motivation
A strong generative model must produce diverse video content and audio types while ensuring fine-grained spatio-temporal synchronization. However, we have noticed that current benchmarks and metrics lack sufficient diversity and robustness for comprehensive evaluation. For instance, the commonly used AIST++ (Li et al., 2021) and Landscape (Lee et al., 2022) benchmarks focus only on limited scenarios (human dancing for AIST++ and natural scenarios for landscape) with merely 20-100 test samples. On the other hand, although TAVGBench (Mao et al., 2024) provides 3,000 test samples (but not released yet), it lacks detailed analysis of different scenarios in practical audio-video generations. While larger-scale datasets such as VGGSound (Chen et al., 2020a) and AudioSet (Gemmeke et al., 2017a) provide millions of video-audio pairs for training and testing, they are typically limited to a single-event taxonomy (e.g., human voice, engine) and do not reflect multi-event dynamics within a single video clip. Consequently, they are not suitable for evaluating spatial-temporal alignment in complex real-world environments. Moreover, all videos in VGGSound are captured in the wild, resulting in a lack of in-house scenes. Both of the two datasets fall short in scene diversity, underrepresenting diverse categories such as 2D/3D animations and industrial settings, which are crucial for evaluating generative models under varied and challenging conditions. In addition, the video-audio synchronization metric AV-Align (Yariv et al., 2024) cannot produce accurate evaluation on complicated scenarios, e.g., with multiple sounding events (Ishii et al., 2024; Mao et al., 2024).
We are therefore motivated to propose a more challenging benchmark to evaluate generative models from various dimensions on a larger data scale in Sec. 4.1, with a particular focus on spatio-temporal synchronization. We also introduce a robust metric to quantify the fine-grained spatio-temporal alignment between generated video and audio pairs in Sec. 4.2. Here we provide more details of the benchmark construction and metric verification.
@@ D.2 Taxonomy
[2pt/2pt] Urban Scenario Outdoor spaces shaped by human activity, including cities, villages, streets, and parks.
[2pt/2pt] Living Scenario Indoor environments where daily human activities occur, like houses, schools, and shopping malls.
[2pt/2pt] Industrial Scenario Work-oriented spaces related to industrial or energy activities, such as factories, construction sites, and mines.
[2pt/2pt] Virtual Scenario Imaginative or abstract settings, including virtual worlds, sci-fi cities, and artistic installations.
@@
[2pt/2pt] 2D-Animate Styles like hand-drawn animation, flat animation, cartoon styles, or watercolor illustrations.
[2pt/2pt] 3D-Animate Photorealistic styles, sci-fi/magical effects, CG (Computer Graphics), or steampunk aesthetics.
@@
[2pt/2pt] Biological Sounds Sounds produced by living creatures (e.g.animals, birds). This includes vocalizations such as barking, chirping, growling, as well as non-vocal human sounds like heartbeat, and other physical noises.
[2pt/2pt] Mechanical Sounds Sounds generated by man-made machines, devices, or mechanical processes. This includes the noise of engines, motors, appliances, and any mechanical or electronic noise. This category also includes malfunction sounds (e.g., malfunctioning machinery or alarms).
[2pt/2pt] Musical Sounds Sounds related to music or musical performance, including both human-generated and instrument-generated sounds and melodies. This category covers singing, instrumental performances, as well as background music used in various media formats.
[2pt/2pt] Speech Sounds Sounds generated from human speech, whether in conversation, dialogue, public speeches, debates, interviews, or monologues. This category specifically covers linguistic communication in various contexts, whether formal, informal, or contentious.
@@
[2pt/2pt] Multiple Subject There are multiple primary objects that (or potentially can) make sounds in the scene.
[2pt/2pt] Off-screen Sound The source of the sound is not visible in the scene but logically exists (e.g., a car engine outside the camera view).
@@
[2pt/2pt] Sequential Events There are multiple events occurring sequentially, with no overlap. For example, “the applause begins after the music performance ends.”
[2pt/2pt] Simultaneous Events Multiple audio sources are present simultaneously, such as “a person speaking while music plays in the background.”
To comprehensively evaluate the capabilities of joint video-audio generation models, we designed five evaluation dimensions from different aspects: (1) Event Scenario, (2) Video Style, (3) Sound Type, (4) Spatial Subjects, and (5) Temporal Composition. Leveraging GPT-4 (Achiam et al., 2023), we enumerated 3-5 primary categories under each aspect, contributing to a hierarchical categorization system with 5 dimensions and 19 categories. Tab. A3 presents the detailed definitions and clarifications.
@@ D.3.1 Collection
We collected data from two sources to construct the benchmark. First, we incorporate existing datasets’ test sets, including those in the JAVG domain like Landscape (Lee et al., 2022) and AIST++ (Li et al., 2021) (despite their limited coverage) and others from sounding-video understanding tasks such as FAVDBench (Shen et al., 2023) and AVSBench (Zhou et al., 2022). Second, we expanded the benchmark by crawling videos uploaded to YouTube between June 2024 and December 2024 to prevent data leakage in previous methods’ training scope (Mao et al., 2024). Using the previously defined categorization system, we prompt GPT4 to generate potential keywords for specific categories, enabling targeted video collection and avoiding indiscriminate gathering of homogeneous data, which significantly improves curation efficiency. This stage yields around 30K sounding video candidates.
@@ D.3.2 Quality-based Pre-Filtering
Following OpenSora (Zheng et al., 2024), we utilize a series of filtering tools to improve data quality, including:
@@
    Aesthetic Filtering. We filter out videos with aesthetic scores (Schuhmann et al., 2021) lower than 4.5, remaining in 70K clips.
@@
    Optical-flow Filtering. We use UniMatch (Xu et al., 2023) to estimate the motion quality of the given videos, and remove the (static) videos whose score is lower than 0.1. This produces 46K candidates.
@@
    OCR Filtering. We use DBNet (Liao et al., 2022) to detect and filter out the videos containing more than 5 text regions, resulting in a smaller 30K scope.
@@ D.3.3 Annotation
  • First, we use Qwen2-VL-72B (Wang et al., 2024c) to generate detailed captions for videos. Metadata from the data source (e.g., object labels or YouTube keywords) is included in the context to enhance caption quality.
  • Next, we use Qwen2-Audio-7B (Chu et al., 2024) to caption the audio in detail. We do not incorporate the previously generated video captions as contextual input, as it produces even more hallucinations with the input bias (i.e., wrongly identifying a sound that will happen in corresponding visual scenarios but actually does not exist in the current audio).
  • Then, we employ Qwen2.5-72B-Instruct (Yang et al., 2024a) to merge video and audio captions into a unified text prompt, serving as the textual condition for the JAVG task. During this process, Qwen2.5-72B is prompted to reason on the video-audio captions and identify apparent logical mistakes (e.g., there is a dog in the video but Qwen2-Audio gives sounds from a car engine) or missing captions (e.g., unknown sound source).
@@ D.3.4 Content-based Post-Filtering
To build a diversified and balanced benchmark to evaluate joint audio-video generation, we conduct another post-filtering based on the categorization results obtained above. In particular, we further remove videos that only contain background music and speech voice. By doing so, nearly half of the videos are filtered out, since a large part of YouTube videos rely on music and voice acting to attract viewers. After human checking, we obtained 10,140 samples in our JavisBench with diverse data sources and fine-grained category annotations, setting a new standard to facilitate a comprehensive evaluation for future joint audio-video generation (JAVG) methods. The category statistics can be found in Fig. 4, and Tab. A4 displays the data sources before and after our filtering strategies.
Table A4: Data source composition before and after filtering strategies of our JavisBench
YouTube FAVDBench AVSBench Landscape AIST++ Total
Before Filtering 30,107 1,000 804 100 20 32,031
After Filtering 8,507 833 680 100 20 10,140
@@ D.4.1 Implementation Details
Given a generated video-audio pair ($V$, AA), we use ImageBind (Girdhar et al., 2023) to estimate audio-visual synchrony with following steps:
@@
    Temporal Segmentation via Sliding Windows. ($V$, AA) is chunked into several segments with 2-seconds window size and 1.5-seconds overlap ($\mathcal{C}=\{(V_{1},A_{1}),(V_{2},A_{2}),\cdots(V_{W},A_{W})\}$). The 2-second window ensures compatibility with the audio encoder’s minimum processing length (Girdhar et al., 2023), preventing suboptimal feature extraction. The 1.5-second overlap enhances continuity and robustness, allowing each frame to be evaluated within multiple temporal contexts. This mitigates artifacts caused by abrupt segmentation boundaries.
@@
    Frame-wise Audio-Visual Similarity Computation. Inspired by Mao et al. (2024), we calculate the cosine similarity between each frames ($\mathcal{F}=\{V_{i,1},V_{i,2},\cdots V_{i,w}\}$) and the whole audio clip $A_{i}$ by using ImageBind’s vision and audio encoders ($E_{v}$, $E_{a}$). Frame-wise similarity captures fine-grained temporal dynamics, ensuring transient events (e.g., lip movements, object interactions) are accurately modeled.
@@
$S_{\textit{Javis}}=\frac{1}{W}\sum_{i=1}^{W}\sigma(V_{i},A_{i}),\quad\sigma(V_{i},A_{i})=\frac{1}{k}\sum_{j=1}^{k}\mathop{\text{top-}k}\limits_{\min}\{\cos\left(E_{v}(V_{i,j}),E_{a}(A_{i})\right)\}\,.$ (A3)